$SPAD/src/input ffrac.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)lib ffrac
f1 : FormalFraction Integer
f1 := 6/3
--       6
--       -
--       3
f2 := (3.6/2.4)$FormalFraction Float
--       3.6
--       ---
--       2.4
numer f1
--       6
denom f2
--       2.4
f1 :: FRAC INT
--       2
% :: FormalFraction Integer
--       2
--       -
--       1
f2 :: Float
--       1.5
output "End of tests"

———- 4

References^{[1] nothing}

5